# Inference Scaling fLaws: The Limits of LLM Resampling with Imperfect Verifiers

### v2

@@ 4 How many samples are optimal?
Figure 4: Inference scaling curves in the presence of a cost for false positives. We show the reward as a function of the number of attempts $K$ across various cost-benefit ratios for the Llama 3.1 and Code Llama model families. Crosses mark the optimal number of samples for each setting. Standard inference scaling curves with no cost (i.e., cost-benefit ratio is 0) are provided for reference. We find that, even at zero computational cost, there is a finite optimal number of samples $K$ that is often very low.
@@
We add one important detail: we model the cost of false positives, such as code that passes unit tests but has subtle bugs. The cost of bugs (which might result in buggy software being deployed) is not easily comparable to the labor-saving benefit of correct solutions, and this cost-benefit ratio can vary greatly depending on the application. So we consider many possible values for the cost-benefit ratio, including zero, which is the setting considered in previous work on inference scaling. The ratio can potentially be much higher than 1 in some applications, such as security sensitive ones, since bugs might translate to exploitable software vulnerabilities.
@@
Experimental setup. For each model of interest, we generated 200 samples for each task in the HumanEval benchmark. For each K $\leq$ 200, If a passing solution was found within $K$ samples, we assigned rewards based on the outcome: a true positive yielded a benefit of 1, while a false positive incurred a cost, with values set according to different cost-benefit ratios: 0, 1, 2, 4, or 8 (Fig. 4). If no passing solution was found within $K$ samples, we assigned a reward of 0 (both, cost and benefit are 0). We repeated this whole process 1,000 times and computed the mean reward for each $K$. The set of samples was the same in all 1,000 runs, but the order of samples was randomly permuted. This setup allows us to empirically observe the relationship between the number of sampling attempts $K$ and the reward for various cost-benefit ratios. The results are illustrated in Fig. 4 for the Llama 3.1 [18] and Code Llama [50] model families. The scaling curves for GPT-4o are included in Fig. 16.
@@
Findings. The results show that the effectiveness of repeated sampling quickly reaches a point of diminishing and even negative returns. Each additional attempt brings a trade-off: although it might yield a correct solution, it might instead yield a false positive, and the false positive rate increases with $K$ (Fig. 5). At first this is surprising, since sampling is a memoryless process. To understand why it happens, we need to look at the distribution of task difficulty (Figure 6), which turns out to be strongly bimodal. The easy tasks get solved within a few attempts, and for the remaining harder tasks, false positives are more likely.
@@
Thus, even with zero computational cost, the optimal number of samples is finite and low (Fig. 4). For example, at a cost-benefit ratio of 4, the optimal number of samples is $K\leq 5$ for all four models. If the ratio is high enough, the optimal number of samples is zero — the expected cost of a false positive outweighs the expected benefit of a correct solution, so the reward is always negative and it is best not to attempt a solution at all.
@@
We note one important caveat: for some models such as Llama 3.1 70B, the false positive rate increases dramatically with $K$, whereas for others such as the Code Llama and Command families, the increase is much more gradual, resulting in much higher values of the optimal $K$, especially for low cost-benefit ratios. We have not been able to identify any intuitive reason for this difference.
@@
To summarize, weaker models cannot “sample their way” to top-level performance if the verifier cannot reliably filter out false positives because the risks quickly outweigh the benefits.
Our findings in this section align with our theoretical model in Section A.3, which generalizes these findings to other benchmarks.
@@
Figure 5: False positive rate as a function of the number of attempts $K$ for Llama 3.1 70B and Code Llama 7B on HumanEval+. We include plots for additional models in Section A.3.
(a) HumanEval+
(b) MBPP+
Figure 6: Distribution of task difficulties for Llama 3.1 8B and GPT-4o on HumanEval+ and MBPP+. We include barplots for all models on both benchmarks in Section A.2.
@@ 5 False positive solutions are low-quality even beyond correctness
While correctness is a fundamental criterion for evaluating code generated by LLMs, it is not the only determinant of code quality. High-quality code possesses attributes beyond mere functionality, such as readability, maintainability, and efficiency. Readability simplifies error-checking and is considered one of the most useful properties of high-quality code [8]. It can be measured using various metrics, including code length guidelines (e.g., PEP8), adherence to naming conventions like snake_case or camelCase, and consistent commenting [79]. Intuitively, shorter code with clear variable names and informative comments is generally easier to read and maintain.
@@
Experimental setup. In our experiments on HumanEval, we evaluate the readability scores of candidate solutions that pass standard unit tests and the more comprehensive test suite. For each measure of code readability, we use a different prompt instructing the model to adhere to the desired guidelines (see Section A.5 for detailed prompts). We rely on the prompts and implementation from Zheng et al. [79].
The results show notable differences in code quality between false positives and robust implementations. False positives, passing only the standard but not the extended unit tests, tend to have worse code quality across all metrics (see Fig. 7). This trend is consistent across models of varying capabilities. This suggests that the limitations of imperfect verifiers for coding tasks extend beyond correctness issues but also affect other code characteristics important for software development. This affects weaker models more, given that they are more prone to generate false positives.
An open question arising from our findings is whether fine-tuning LLMs on code quality metrics could improve not only the quality of generated code but also robustness [29], potentially mitigating the prevalence of false positives.
Figure 7: False positives tend to be lower-quality code than correct implementations. For example, false positive solutions fail to adhere to the camelCase naming convention more often than robust implementations. Fig. 17 shows that this holds consistently across models and for all four code quality metrics we test. GPT-3.5 exhibits a low relative performance in using camelCase but performs comparably to other models in following snake_case(Section A.5). This has also been found in previous work [79].
Qualitative analysis of false positives. To better understand the characteristics of false positive implementations, we randomly sample 10 implementations across all models that pass the standard tests but fail the additional unit tests, 5 for each benchmark. Through manual analysis, we identified several recurring error types. All examples mentioned in the following are included in Section A.5.3.
@@
    Logical errors: Such errors were common. For instance, in the HumanEval/30 task, the model is tasked with returning only positive numbers from a list. The generated solution shown in Fig. 8(b) incorrectly converts positive floats to integers, passing the basic tests that included only integers but failing extended tests that introduced float values.
@@
    Inefficient implementations: While most false positives result from logic errors or edge case mishandling, some were also caused by inefficient implementations. For instance, in HumanEval/15, the solution involved a for-loop that became inefficient when handling larger inputs, causing a timeout on the extended tests. Following the implementation of Liu et al. [42], we set the timeout such that each candidate solution must compute in less than one second or four times the time it takes to run each test on the ground truth implementation, whichever is greater.
1def get_positive(l: list):
2 return [int(item) for item in l if item > 0]
(a) Example implementation generated by CodeGen-6B which fails additional unit tests from HumanEval+.
1# Standard unit tests
3assert get_positive([-1, 2, -4, 5, 6]) == [2, 5, 6]
4
5# Example unit test from extended test suite
(b) Standard unit tests for the HumanEval/30 task and one example test from the extended test suite of HumanEval+.
Figure 8: Example implementation that passes the standard tests but is incorrect. In HumanEval/30, the model is asked to Return only positive numbers in the list. The code was generated by CodeGen-6B [48] and passes the tests provided with the original HumanEval benchmark but fails the extended tests. The implementation contains a logical error due to unnecessarily converting positive floats to integers.
@@ 6 Discussion
We study a setting where all generators are paired with the same verifier. The verifier has imperfect coverage, but no false negatives (i.e. it doesn’t reject correct solutions). In real-world deployment settings, human-written unit tests are rarely available and we would need to rely on the use of automated test generation techniques. These approaches include symbolic execution [46], specialized transformers [62], and LLMs [12, 9, 56]. Model-generated tests introduce new challenges including a disparity between verifiers and a risk of false negatives. This could widen the generalization gap. Recent work has sought to improve test quality by increasing coverage [52, 1], identifying edge cases [46], or reproducing user-specific bugs [31, 43]. We leave an investigation of the impact of model-generated unit tests as a next step.
Our findings weaken the support for our previous paper’s claim that evaluations must be cost-controlled in order to be useful [32]; here, even when disregarding computational cost, resampling with imperfect verifiers is inherently limited.
@@
Limitations. Our experiments focus solely on repeated sampling in the context of coding tasks. While this domain offers a clear example of the challenges posed by imperfect verifiers, other domains might exhibit different behavior. Future work could extend these findings to tasks such as reasoning [25], web agents [5, 23], or agent-user interaction [76]. Another limitation is prompt sensitivity, which affects LLM evaluations [6, 38]. While we followed the original authors’ implementation provided with the HumanEval+ and MBPP+ benchmarks [42], prompt engineering could influence false positive generation. Additionally, we did not investigate how benchmark contamination contributes to our findings, as models could be overly optimized for passing the standard test cases. Finally, we did not explore mitigation strategies such as training models to improve robustness against verifier weaknesses or refining solutions after they passed the verifier [51]. Similarly, we did not test alternative strategies to inference scaling that, e.g., induce more diversity during sampling, such as PlanSearch [65].
Acknowledgments
@@
We thank Boris Hanin, Peter Henderson, Alexander Wettig, Jon Saad-Falcon, and Zachary Siegel for providing valuable feedback on earlier drafts of this work.
Appendix A Appendix
A.1 Additional details on Section 2
A.1.1 Edge cases in our generator-verifier setting
@@
The setting described in Fig. 2 considers verifiers and generators as distinct components, where verifiers score and accept or reject individual samples from the generator’s output to enable accuracy improvements through resampling.
@@
This creates interesting edge cases with methods like Chain-of-Verification (CoVe) [17], Tree of Thoughts (ToT) [75], and Reward Models (RMs) [57] where verification and generation are more tightly coupled. While ToT fits within our framework by producing aggregate scores and a decision on whether the problem is solvable from a given state or not (i.e. potentially rejecting solutions), CoVe differs fundamentally in its verification approach. Instead of producing numeric scores and accepting or rejecting solution candidates, CoVe uses verification of intermediate facts used for answering a question to improve a single response through iterative refinement. This makes CoVe less suitable for inference scaling through resampling because there is no way to distinguish between the quality of multiple samples and using a verifier’s verdict for resampling.
A.2 Additional details on Section 3
@@
(a) HumanEval+
(b) MBPP+
Figure 13: Upper and lower bounds for our HumanEval+ and MBPP+ analysis. We show the relationship between the accuracy of individual samples (x-axis) and the achievable accuracy given an infinite compute budget and limited unit tests (y-axis; note that it starts at 0.5). Error bars indicate the bounds on the conditional accuracy when we account for tasks for which we did not observe any passing solutions. For the upper (lower) bound, we set the conditional passing rate to 1 (0) when computing the accuracy estimates. Note that for Command-Light, even after collecting 1000 samples for each task on HumanEval+, we still observed a substantial fraction of the tasks without a single passing solution.
A.2.1 Details on data and implementation
@@
Sample Collection. To evaluate the generalization gap between weaker and stronger models, we collected multiple samples per model and task. For both benchmarks and each model, we used samples generated with a temperature setting of 0.8. For sample generation, we use the implementation provided by Liu et al. [42], and other than the temperature use their default settings. [^1] See: https://github.com/evalplus/evalplus/tree/937c46858cf8e687b31b5a728b7083d6e5a84971 We collected a minimum of 50 samples for each model and task. For most models in our experiments Vicuna 7B, Mistral 7B, CodeT5p 16B, CodeGen, CodeGen2, Code Llama 7B, and Code Llama 13B, we used samples made available by Liu et al. [42]. These were collected using the same temperature setting (i.e., 0.8). We had access to 200 samples per model and task for these models. Additionally, on HumanEval+, we collected 200 samples for Llama 3.1, Phi-3, GPT-4o and the Command family of models. For Command-Light, we even collected 1000 samples for each task to reduce the number of tasks without any solutions passing the HumanEval unit tests (Fig. 13). On MBPP+, we collected 50 samples for each model and task.
A.2.2 Additional details on excluded tasks from MBPP+

### v3

@@ 4 How many samples are optimal?
We add one important detail: we model the cost of false positives, such as code that passes unit tests but has subtle bugs. The cost of bugs (which might result in buggy software being deployed) is not easily comparable to the labor-saving benefit of correct solutions, and this cost-benefit ratio can vary greatly depending on the application. So we consider many possible values for the cost-benefit ratio, including zero, which is the setting considered in previous work on inference scaling. The ratio can potentially be much higher than 1 in some applications, such as security sensitive ones, since bugs might translate to exploitable vulnerabilities.
@@
Experimental setup. For each model of interest, we generated 200 samples for each task in the HumanEval benchmark. For each K $\leq$ 200, If a passing solution was found within $K$ samples, we assigned rewards based on the outcome: a true positive yielded a benefit of 1, while a false positive incurred a cost, with values set according to different cost-benefit ratios: 0, 1, 2, 4, or 8 (Figure˜4). If no passing solution was found within $K$ samples, we assigned a reward of 0 (both, cost and benefit are 0). We repeated this whole process 1,000 times and computed the mean reward for each $K$. The set of samples was the same in all 1,000 runs, but the order of samples was randomly permuted. This setup allows us to empirically observe the relationship between the number of sampling attempts $K$ and the reward for various cost-benefit ratios. The results are illustrated in Figure˜4 for the Llama-3.1 (Dubey et al., 2024) and Code Llama (Rozière et al., 2024) model families. The scaling curves for GPT-4o are included in Figure˜15.
Figure 4: Inference scaling curves in the presence of a cost for false positives. We show the reward as a function of the number of attempts $K$ across various cost-benefit ratios for the Llama-3.1 and Code Llama model families. Crosses mark the optimal number of samples for each setting. Standard inference scaling curves with no cost (i.e., cost-benefit ratio is 0) are provided for reference. We find that, even at zero computational cost, there is a finite optimal number of samples $K$ that is often very low.
@@
Findings. The results show that the effectiveness of repeated sampling quickly reaches a point of diminishing and even negative returns. Each additional attempt brings a trade-off: although it might yield a correct solution, it might instead yield a false positive, and the false positive rate increases with $K$ (Figure˜5). At first this is surprising, since sampling is a memoryless process. To understand why it happens, we need to look at the distribution of task difficulty (Figure 6), which turns out to be strongly bimodal. The easy tasks get solved within a few attempts, and for the remaining harder tasks, false positives are more likely. This aligns with findings by Chen et al. (2024a), who observed a similar inverse U-shaped accuracy curves explained by the heterogeneity in task difficulties.
@@
Thus, even with zero computational cost, the optimal number of samples is finite and low (Figure˜4). For example, at a cost-benefit ratio of 4, the optimal number of samples is $K\leq 5$ for all four models. If the ratio is high enough, the optimal number of samples is zero — the expected cost of a false positive outweighs the expected benefit of a correct solution, so the reward is always negative and it is best not to attempt a solution at all.
@@
We note one important caveat: for some models such as Llama-3.1-70B, the false positive rate increases dramatically with $K$, whereas for others such as the Code Llama and Command families, the increase is much more gradual, resulting in much higher values of the optimal $K$, especially for low cost-benefit ratios. We have not been able to identify any intuitive reason for this difference.
@@
To summarize, weaker models cannot “sample their way” to top-level performance if the verifier cannot reliably filter out false positives because the risks quickly outweigh the benefits. Our findings in this section align with our theoretical model in Appendix˜C, which generalizes these findings to other benchmarks.
@@
Figure 5: False positive rate as a function of the number of attempts $K$ for Llama 3.1 70B and Code Llama 7B on HumanEval+. We include plots for additional models in Appendix˜C.
@@ 5 False positive solutions are low-quality even beyond correctness
While correctness is a fundamental criterion for evaluating code generated by LLMs, it is not the only determinant of code quality. High-quality code possesses attributes beyond mere functionality, such as readability, maintainability, and efficiency. Readability simplifies error-checking and is considered one of the most useful properties of high-quality code (Börstler et al., 2023). It can be measured using various metrics, including code length guidelines (e.g., PEP8), adherence to naming conventions like snake_case or camelCase, and consistent commenting (Zheng et al., 2024). Intuitively, shorter code with clear variable names and informative comments is generally easier to read and maintain.
@@
Experimental setup. On HumanEval, we evaluate the readability scores of candidate solutions that pass standard unit tests and the more comprehensive test suite. For each measure of code readability, we use a different prompt instructing the model to adhere to the desired guidelines (see Appendix˜D for detailed prompts). We rely on the prompts and implementation from Zheng et al. (2024).
The results show notable differences in code quality between false positives and robust implementations. False positives, passing only the standard but not the extended unit tests, tend to have worse code quality across all metrics (see Figure˜7). This trend is consistent across models of varying capabilities. This suggests that the limitations of imperfect verifiers for coding tasks extend beyond correctness issues but also affect other code characteristics important for software development. This affects weaker models more, given that they are more prone to generate false positives.
An open question arising from our findings is whether fine-tuning LLMs on code quality metrics could improve not only the quality of generated code but also robustness (Jain et al., 2023), potentially mitigating the prevalence of false positives.
HumanEval+
MBPP+
Figure 6: Distribution of task difficulties for Llama 3.1 8B and GPT-4o on HumanEval+ and MBPP+. We include barplots for all models on both benchmarks in Appendix˜B.
Figure 7: False positives tend to be lower-quality code than correct implementations. For example, false positive solutions fail to adhere to the camelCase naming convention more often than robust implementations. Figure˜16 shows that this holds consistently across models and for all four code quality metrics we test. GPT-3.5 exhibits a low relative performance in using camelCase but performs comparably to other models in following snake_case (Appendix˜D). This has also been found in previous work (Zheng et al., 2024).
Qualitative analysis of false positives. To better understand the characteristics of false positive implementations, we randomly sample 10 implementations across all models that pass the standard tests but fail the additional unit tests, 5 for each benchmark. Through manual analysis, we identified several recurring error types. All examples mentioned in the following are included in Section˜D.3.
@@
    Logical errors: Such errors were common. For instance, in the HumanEval/30 task, the model is tasked with returning only positive numbers from a list. The solution shown in Figure˜18 incorrectly converts floats to integers, passing the basic tests that included only integers but failing extended tests that introduce float values.
@@
    Inefficient implementations: While most false positives result from logic errors or edge case mishandling, some were also caused by inefficient implementations. For instance, in HumanEval/15, the solution involved a for-loop that became inefficient when handling larger inputs, causing a timeout on the extended tests. Following Liu et al. (2023b), we set the timeout such that each candidate solution must compute in less than one second or four times the time it takes to run each test on the ground truth implementation, whichever is greater.
@@ 6 Discussion
We study a setting where all generators are paired with the same verifier. The verifier has imperfect coverage, but no false negatives. In real-world deployment settings, human-written unit tests are rarely available and we would need to rely on the use of automated test generation techniques. These approaches include symbolic execution (Lukasczyk and Fraser, 2022), specialized transformers (Tufano et al., 2021), and LLMs (Chen et al., 2024b; 2022; Siddiq et al., 2024). Model-generated tests introduce new challenges including a disparity between verifiers and a risk of false negatives. This could widen the generalization gap. We leave an investigation of the impact of model-generated unit tests as a next step.
Resampling is used not only to scale inference but also to train large reasoning models. Many state-of-the-art reasoning models are trained on datasets curated through rejection sampling (NovaSky Team, 2025; DeepSeek-AI et al., 2025; Bespoke Labs, 2025), where verifiers filter out incorrect outputs. However, imperfect verifiers can introduce mislabeled examples, implicitly incurring a cost of false positives. We hypothesize that weaker models paired with imperfect verifiers fail to produce datasets of sufficient quality to train competitive reasoning models, creating a bottleneck: without stronger base models or more accurate verifiers, gains from resampling-based data curation to train reasoning models are limited.
Although our experiments focus on coding tasks, we want to emphasize that our theoretical results are domain-agnostic: resampling relying on imperfect verifiers with non-zero false positive rates will face the same fundamental ceiling.
Related work has also studied the risks of over-optimizing against imperfect rewards. For example, Gao et al. (2022) analyzes how optimizing against proxy reward models can lead to degraded true performance. Our setting differs in that we focus on inference-time resampling, where the ceiling on achievable accuracy arises directly from false positives rather than reward misalignment.
Finally, our findings weaken support for previous papers’ claims that resampling is an effective strategy to increase accuracy by trading off inference time compute (Kapoor et al., 2024; Chen et al., 2024a); here resampling with imperfect verifiers is inherently limited.
@@
Limitations. Our experiments focus on repeated sampling in the context of coding tasks. Coding offers a clear example of the challenges posed by imperfect verifiers, other domains might exhibit different behavior. Future work could extend these findings to tasks such as reasoning (Hosseini et al., 2024a), web agents (Bai et al., 2024; He et al., 2024), or agent-user interaction (Yao et al., 2024). Another limitation is prompt sensitivity, which affects LLM evaluations (Biderman et al., 2024; Liang et al., 2023). While we followed the original authors’ implementation provided with the HumanEval+ and MBPP+ benchmarks (Liu et al., 2023b), prompt engineering could influence false positive generation. Additionally, we did not investigate how benchmark contamination affects our findings, as models could be overly optimized for passing the standard test cases. We did not explore mitigation strategies such refining solutions after they passed the verifier (Saad-Falcon et al., 2024). Similarly, we did not test alternative strategies to inference scaling that, e.g., induce more diversity during sampling (Wang et al., 2024a). Finally, the optimal number of samples K that we identify depends on the cost-benefit ratio, which varies by application context. While we show that K is finite and often very low across a range of ratios, we do not provide empirical evidence for what specific ratio applies in any given real-world deployment. In security-critical applications, the cost of a false positive may far exceed the benefit of a correct solution, whereas in lower-stakes settings a higher K may remain practical.
@@
Acknowledgments. We thank Boris Hanin, Peter Henderson, Alexander Wettig, Jon Saad-Falcon, and Zachary Siegel for providing valuable feedback on earlier drafts of this work. We thank Cohere for their support of this work with API credits. We also thank Schmidt Sciences for their funding and support of this work.
Reproducibility. We release code to reproduce all experimental results of this paper in a GitHub repository [^1] URL: https://github.com/benediktstroebl/inference-scaling-limits. This repository also contains all code samples for all models used in our experiments. Additionally, we provide an implementation of the theoretical model in Appendix˜C as a Python notebook.
Appendix A Additional details on Section˜2
A.1 Edge cases in our generator-verifier setting
@@
The setting described in Figure˜2 considers verifiers and generators as distinct components, where verifiers score and accept or reject individual samples from the generator’s output to enable accuracy improvements through resampling.
@@
This creates interesting edge cases with methods like Chain-of-Verification (CoVe) (Dhuliawala et al., 2023), Tree of Thoughts (ToT) (Yao et al., 2023), and Reward Models (RMs) (Snell et al., 2024) where verification and generation are more tightly coupled. While ToT fits within our framework by producing aggregate scores and a decision on whether the problem is solvable from a given state or not (i.e. potentially rejecting solutions), CoVe differs fundamentally in its verification approach. Instead of producing numeric scores and accepting or rejecting solution candidates, CoVe uses verification of intermediate facts used for answering a question to improve a single response through iterative refinement. This makes CoVe less suitable for inference scaling through resampling because there is no way to distinguish between the quality of multiple samples and using a verifier’s verdict for resampling.
Appendix B Additional details on Section˜3
@@
HumanEval+
MBPP+
Figure 12: Worst-case upper and lower bounds for our HumanEval+ and MBPP+ analysis. We show the relationship between the accuracy of individual samples (x-axis) and the achievable accuracy given an infinite compute budget and limited unit tests (y-axis; note that it starts at 0.5). Error bars indicate the bounds on the conditional accuracy when we account for tasks for which we did not observe any passing solutions. For the upper (lower) bound, we set the conditional passing rate to 1 (0) when computing the accuracy estimates. Note that for Command Light, even after collecting 1000 samples for each task on HumanEval+, we still observed a substantial fraction of the tasks without a single passing solution.
B.1 Details on data and implementation
@@
Sample Collection. To evaluate the generalization gap between weaker and stronger models, we collected multiple samples per model and task. For both benchmarks and each model, we used samples generated with a temperature setting of 0.8. For sample generation, we use the implementation provided by Liu et al. (2023b), and other than the temperature use their default settings. [^2] See: https://github.com/evalplus/evalplus/tree/937c46858cf8e687b31b5a728b7083d6e5a84971 We collected a minimum of 50 samples for each model and task. For most models in our experiments Vicuna 7B, Mistral 7B, CodeT5p 16B, CodeGen, CodeGen2, Code Llama 7B, and Code Llama 13B, we used samples made available by Liu et al. (2023b). These were collected using the same temperature setting (i.e., 0.8). We had access to 200 samples per model and task for these models. Additionally, on HumanEval+, we collected 200 samples for Llama 3.1, Phi-3, GPT-4o, and the Command family of models. For Command Light, we even collected 1000 samples for each task to reduce the number of tasks without any solutions passing the HumanEval unit tests (Figure˜12). On MBPP+, we collected 50 samples for each model and task.
B.2 Additional details on excluded tasks from MBPP+